I need help on a couple questions? These are multiple choice.
A basketball drops from a height of 76 in. On bounce 1 it rebounds to a height of 38 in. On bounce 2 it rebounds to a height of 19 in. How high does the basketball bounce on bounce 3?
a. 9 in. B. 19 in. C. 9.5 in. D 4.75 in.

Write a rule for the pattern: -3,-9,-14, -42,-47,-141.....?????
Please HELP!!!!!

Answer by oscargut(2103)

it rebounds always the half
76,38,19 so the answer is 9.5
-3,-9,-14, -42,-47,-141
first you do -3*3=-9 after that you substract 5
-9-5=-14
after that you do -14*3=-42
the pattern is :
multply by 3, substract 5.......
-3*3=-9
-9-5=-14
-14*3=-42
-42-5=-47
-47*3=-141
-141-5=-146
-146*3=-438
......